Solve: 6sin square x - sinxcosx - cos square x =3

Solve:


6sin square x - sinxcosx - cos square x =3

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

use cos^2(x) =(1+cos2x)/2 sin^2(x) =(1-cos2x)/2 and sinxcosx = sin(2x)/2, solve the resulting equation for 2x and then divide by 2 to get x

vijesh jayan
12 years ago

6sin2x-sinx.cosx-cos2x = 3

multiplying 2 on both sides we get

12sin2x-2sinx.cosx-2cos2x = 6

but we know that cos2x=1-sin2x  and  2sinx.cosx=sin2x

12sin2x-sin2x-2(1-sin2x) = 6

12sin2x-sin2x-2+2sin2x = 6

14sin2x-sin2x = 6+2 = 8

further simplify